### v1

# Beyond the Answers: Reviewing the Rationality of Multiple Choice Question Answering for the Evaluation of Large Language Models

Abstract
In the field of natural language processing (NLP), Large Language Models (LLMs) have precipitated a paradigm shift, markedly enhancing performance in natural language generation tasks. Despite these advancements, the comprehensive evaluation of LLMs remains an inevitable challenge for the community. Recently, the utilization of Multiple Choice Question Answering (MCQA) as a benchmark for LLMs has gained considerable traction. This study investigates the rationality of MCQA as an evaluation method for LLMs. If LLMs genuinely understand the semantics of questions, their performance should exhibit consistency across the varied configurations derived from the same questions. Contrary to this expectation, our empirical findings suggest a notable disparity in the consistency of LLM responses, which we define as REsponse VAriability Syndrome (REVAS) of the LLMs, indicating that current MCQA-based benchmarks may not adequately capture the true capabilities of LLMs, which underscores the need for more robust evaluation mechanisms in assessing the performance of LLMs.
Machine Learning, ICML
Research Center for Social Computing and Information Retrieval, Harbin Institute of Technology, China
{hcwang,sdzhao@ir.hit.edu.cn}
1 Introduction
The emergence of Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023a), Baichuan (Yang et al., 2023), and ChatGPT (OpenAI, 2022), represents a paradigm shift in the field of Natural Language Processing (NLP). These models have exhibited exceptional proficiency in mimicking human-like textual outputs, establishing their significance across a diverse array of applications. However, the challenge of effectively evaluating LLMs persists (Chang et al., 2023). This difficulty arises from the intricate nature of natural language. Conventional evaluation metrics for generative tasks often fall short in accurately assessing the performance of LLMs, since most LLMs can generate text contextually rich and coherent (Thoppilan et al., 2022), complicating the assessment of the outputs through merely quantitative evaluation based on text matching such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004).
Multiple-Choice Question Answering (MCQA) is a fundamental format for various tasks in NLP, such as common sense reasoning (Talmor et al., 2019; Sap et al., 2019; Zellers et al., 2018), reading comprehension (Lai et al., 2017; Huang et al., 2019) and cloze-style tasks (Zellers et al., 2019; Mostafazadeh et al., 2016). Each MCQA instance comprises a question paired with several answer options, requiring models to identify the correct response as depicted in Figure 1. As a non-subjective metric, MCQA serves as a prominent automatic evaluation method with accuracy as evaluation metric for numerous LLMs to test for the commonsense knowledge or knowledge for specific domain (Gao et al., 2021; Touvron et al., 2023a; OpenAI et al., 2023).
Figure 1: An example of Multiple-Choice Question Answering and two ranking strategies.
Despite the advanced performance of LLMs on the accuracy of MCQA-format benchmarks like MMLU (Hendrycks et al., 2021), a key challenge that persists in evaluating LLMs is maintaining consistency in responses when confronted with variations in answer options, and discrepancies arise even when the same questions are reformulated into the others (Robinson & Wingate, 2022; Wang et al., 2023; Zheng et al., 2023). The prior studies have revealed that LLMs exhibit sensitivity to the order of answer choices, leading to inconsistent responses, which underscores an issue that the accuracy of MCQA-format tasks may not reflect the authentic capability of LLMs. However, the fluctuation in the LLM responses due to the order of answer options may not be the only hurdle in the evaluation of LLMs.
In this study, we undertake a comprehensive examination through experiments involving various configurations based on the MCQA-format datasets: (1) original MCQA datasets, (2) modified MCQA datasets with altered answer options, and (3) conversion of MCQA-format instances into true/false format. In conducting this comprehensive evaluation, we deepen the insight into the limitations of LLMs in terms of response consistency. We define the phenomenon as REsponse VAriability Syndrome (REVAS), where the LLMs show inconsistencies in the questions derived from one question. Empirical results have proved that the high performance of the MCQA-format benchmark does not denote that the LLMs fully understand the semantics of the questions or learn the knowledge about the questions. Additionally, we explore the underlying reasons for observed behavioral patterns in LLMs through substantial experiments, that LLMs probably choose the most correct from the answer options and meanwhile regard some other options as correct as well. Therefore, we propose complementary training with negative samples to refine the LLMs, which could pave the way for developing more robust and adaptable NLP models, ultimately bridging the gap between machine and human-like understanding and reasoning in NLP tasks.
In summary, our contributions are as follows:
  • We challenge the rationality of evaluating LLMs using MCQA-format benchmarks by involving various instances derived from the MCQA tasks.
  • Empirical findings suggest that LLMs may acquire a limited understanding of the MCQA-format questions, exhibiting limited generalization even with minor variations.
  • We hypothesize underlying causes for the observed patterns in LLMs and suggest potential strategies for enhancing the robustness and adaptability of LLMs.
2 Related Work
Transformer-based language models (Vaswani et al., 2017) have led the research of nature language processing (NLP) into a new era with the structure of self-attention and massive pre-training on large amounts of self-supervised corpora. Prior work has elucidated the scaling law, positing that the efficacy of language models is typically proportional to both the model size and the magnitude of pre-training data (Henighan et al., 2020; Kaplan et al., 2020). This principle has catalyzed a paradigm shift towards Large Language Models (LLMs), with the number of parameters escalating from millions (Devlin et al., 2019; Radford et al., 2019) to billions (Radford et al., 2018; Brown et al., 2020).
Recent advancements, including supervised fine-tuning and alignment with human values (Ouyang et al., 2022; Chung et al., 2022; Bai et al., 2022), have further augmented the capabilities of LLMs, enabling them to adhere more closely to human instructions and ethical considerations. Nonetheless, challenges persist since LLMs may show inconsistency in the model responses, especially under the scenarios of MCQA. Robinson & Wingate 2022 termed the ability to associate the answer options and corresponding symbols as multiple choice symbol binding (MCSB) and proved that the MCSB ability varied significantly by models. Additionally, Wang et al. 2023 revealed vulnerabilities in the ranking of candidate responses, which could be manipulated by altering the presentation order. Zheng et al. 2023 investigated the token bias in LLMs, which could lead to selection biases during option prediction. Furthermore, Kadavath et al. 2022 explored the reliability of the LLM performance and calibration, focusing exclusively on a set of private models under the MCQA settings. Their methodology involved modifying the options and transforming the original questions into binary-choice questions answering with options of “True” and “False”. Building on these foundational studies, in this study we conduct a comprehensive range of experiments applied to a diverse series of LLMs, including both open-source and commercial models. We aim to assess their performance consistency across various settings while using the same set of instances, thereby providing a broader understanding of their capabilities and limitations in different application scenarios.
Table 1: Common LLMs and corresponding evaluation tasks involving MCQA.
Chinchilla Falcon FLAN Galactica GPT4 LLaMA2 Mistral PaLM2
MMLU
Big Bench
CommonsenseQA
AGI Eval
3 Can the Accuracy of MCQA-format Tasks Reflect the True Model Capability?
3.1 Overview
The assessment of LLMs has increasingly employed MCQA as a standardized metric. This method, featuring a single question accompanied by multiple answer choices, serves to gauge the performance of an LLM performance in specific tasks or domains. Notable MCQA datasets utilized in evaluating LLMs encompass a diverse range including multitask language comprehension (e.g., MMLU (Hendrycks et al., 2021)), benchmark assessments for LLMs (e.g., Big Bench (Srivastava et al., 2023)), commonsense reasoning (e.g., HellaSwag (Zellers et al., 2019), ARC (Clark et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021)), and human-centric evaluations (e.g., AGI-eval (Zhong et al., 2023)). Table 1 lists prominent LLMs such as Chinchilla (Hoffmann et al., 2022), Falcon (Almazrouei et al., 2023), FLAN (Chung et al., 2022), Galactica (Taylor et al., 2022), GPT4 (OpenAI et al., 2023), LLaMA2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and PaLM2 (Anil et al., 2023), along with their respective evaluations with MCQA-format tasks. Additionally, MCQA tasks extend to tests like AP exams for GPT4 and medical-specific MultiMedQA for Med-PaLM (Singhal et al., 2023). The adoption of MCQA tasks as evaluation of LLMs is based on the hypothesis that MCQA can effectively and accurately reveal the true capabilities of the LLMs. If so, LLMs must demonstrate an understanding of the contextual content and possess the requisite knowledge for correctly answering questions, thereby exhibiting consistencies across similar questions. To examine the above hypothesis, we conduct a series of experiments using configurations derived from several MCQA datasets, aimed at observing the behavior of the LLMs.
3.2 Experimental Design
In our study, we select three widely recognized LLMs: LLaMA2 (Touvron et al., 2023b), Baichuan2 (Yang et al., 2023), and ChatGPT 3.5 (OpenAI, 2022), for evaluation on the MMLU (Hendrycks et al., 2021) benchmark and the MedMCQA (Pal et al., 2022) dataset. MMLU represents a prevalent benchmark in MCQA for LLMs as presented in Table 1, while MedMCQA demands extensive medical knowledge and is challenging for most LLMs. The selected evaluation samples comprise 452 instances from MMLU and 967 instances from MedMCQA. These instances are specifically chosen for their adaptability to alternative formats, allowing for the creation of derived datasets. These derived datasets, encompassing a magnified sample size—37 times that of the initially selected instances—facilitate a more extensive and nuanced analysis of LLM capabilities, culminating in over 52,000 testing instances. Further details regarding the selected and constructed instances are available in the supplementary materials. Each experiment is conducted in both zero-shot and few-shot settings.
3.3 Adaptation of MCQA Answer Options
Figure 2: Experimental designs derived from original answer options in the format of MCQA.
Figure 3: Performance of LLMs on MCQA tasks of MMLU and MedMCQA with a baseline of 4 (original) along with variations in the number of options. Accuracy is reported with the initial answer option order. Sc denotes the mean value of $S_{c}$. Acc-H: accuracy-hard, that the prediction of an instance is deemed correct only when LLMs predict consistently and correctly across all permutations of re-ordered options for a given instance. Marked values in the sub-figures are with 4 original options under the 0-shot setting.
Our primary method involves subtle modifications to the answer options within MCQA instances (as depicted in Figure 2). Ideally, an intelligent LLM should display resilience to such modifications, since the knowledge required by the question remains unchanged.
Table 2: Distribution of $S_{c}$ (and corresponding accuracy) on MMLU with the original answer options under the 0-shot setting. %$S_{c}=n/4$ indicates the proportion of instances where LLMs consistently predicted $n$ out of 4 questions with re-ordered options. N/A: not applicable.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
%$S_{c}=4/4$ 43.02 (62.74) 19.03 (67.39) 48.23 (81.65) 34.54 (61.38) 52.43 (85.23) 40.43 (85.17)
%$S_{c}=3/4$ 34.73 (77.71) 37.83 (64.33) 38.27 (44.51) 40.74 (42.13) 35.17 (40.93) 39.81 (43.48)
%$S_{c}=2/4$ 20.35 (21.02) 29.37 (56.52) 11.28 (35.29) 17.89 (31.79) 11.50 (7.59) 17.89 (14.07)
%$S_{c}=1/4$ 7.08 (N/A) 8.58 (N/A) 2.21 (N/A) 6.83 (N/A) 0.88 (N/A) 1.86 (N/A)
Table 3: Performance deterioration in the LLMs post “None of the Above” substitution. Acc denotes the accuracy of LLMs with the initial order of the other options. Acc-H, termed as accuracy-hard, implies that an instance is deemed correctly predicted only if the LLMs demonstrate consistency across every permutation of the answer options. Red font delineates the performance decrement compared to that with the original options.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
0-shot Acc 16.59 (-56.14) 20.27 (-22.96) 11.73 (-59.48) 14.99 (-30.82) 15.93 (-49.07) 27.09 (-29.79)
Acc-H 0.22 (-26.77) 1.34 (-11.48) 3.32 (-26.06) 7.24 (-13.96) 4.65 (-40.04) 8.07 (-26.37)
5-shot Acc 3.54 (-70.90) 4.45 (-38.88) 11.73 (-59.53) 14.99 (-28.96) 34.73 (-55.07) 40.85 (-18.92)
Acc-H 0.44 (-30.53) 0.31 (-12.62) 1.11 (-28.53) 1.76 (-12.20) 9.29 (-34.07) 15.51 (-16.46)
Re-ordering of Answer Options
Previous studies, such as Robinson & Wingate 2022, have demonstrated that re-ordering answer options can induce performance variability, a phenomenon intrinsically linked to the nature of LLMs. For simplicity, we expand each chosen instance into four random distinct orders to assess the sensitivity of LLMs to option sequencing. We conduct experiments with MCQA in its original format on selected instances from MMLU and MedMCQA, recording baseline accuracy (“Acc”) just like the reported accuracy for evaluating LLMs. Subsequently, we introduce a consistency score “$S_{c}$” defined for each instance as follows:
$S_{c}=\frac{\texttt{UNIQUE\_ANSWER}\ (a_{1},\ \dots,\ a_{n})}{n}$
$\bar{S}_{c}=\frac{1}{N}\sum_{i=1}^{N}{S_{c}(i)}$
where $n$ represents the number of options in the MCQA task, and $N$ is the total number of task instances. For MCQA tasks with $n$ options, $S_{c}$ ranges from $1/n$ to $1.0$ describing the proportion of consistent answers with $n$ orders, where a score of 1.0 signifies consistent LLM performance across all re-ordered questions. Additionally, we propose a novel evaluation metric, termed accuracy-hard (“Acc-H”), which measures the proportion of instances where LLMs consistently and correctly predict across all re-ordered variations.
Variation in Answer Option Quantity
In addition to re-ordered options, our investigation extends to evaluating the influence of varying the number of answer choices. Since the norm of four or five options in most MCQA assessments (Haladyna & Downing, 1993), we deliberate modifications in the number of answer options to examine the adaptability of LLMs to these altered conditions.
When incrementing the number of answer options, we incorporate additional non-semantic tokens as distracting options to circumvent the inadvertent introduction of plausible correct responses, thereby preserving the authenticity of the questions. This expands the number of answer options to 6, 8, and 10 choices. In contrast, when reducing the number of answer options, we randomly eliminate distracting options, simplifying the questions by reducing the options to 3 or 2.
By combining the adaptions of re-ordering and quantity variation of options, Figure 3 demonstrates the performance of LLMs on the MMLU and MedMCQA tasks. The metric “Accuracy” is quantified based on questions with only a single option order for $n$ options. The $\bar{S}_{c}$ and Acc-H represent the mean values quantified across $n$ option orders for $n$ options respectively. Illustratively, taking Baichuan2-13B as an example, it exhibits an accuracy of 70.21% with the original set of four options on MMLU under the zero-shot setting, alongside a notable mean $S_{c}$ of 83.13. However, the Acc-H metric indicates a significant drop, with consistent and correct predictions in merely 39.38% of all instances when each instance is re-ordered into four distinct instances. Observing the performance of Baichuan2-13B with 2 to 10 answer options, a marked volatility in Accuracy is noted, ranging from 44.44% to 75.45%, with the peak accuracy surprisingly achieved with 10 options. Nonetheless, both $S_{c}$ and Acc-H exhibit a declining trend as the number of options increases. This pattern of fluctuating accuracy and diminishing $S_{c}$ and Acc-H values is consistent across all three LLMs on the MMLU. In experiments involving MedMCQA, a similar downward trend is observed for Accuracy, $S_{c}$, and Acc-H with the increment in answer options. Remarkably, for LLaMA2-13B, Acc-H plummets to 0.00% with six or more answer options, casting doubts on the reliability of the LLMs with a larger array of answer options. These observations underscore the questionable reliability of Accuracy metrics when based on a single order of answer options.
For a fine-grained analysis, Table 2 presents the distribution of $S_{c}$ values under the zero-shot setting. It is noticeable that most of the LLMs achieve an $S_{c}$ of at least 3/4. Notably, a higher $S_{c}$ value correlates with higher accuracy in the LLM responses. For instance, the accuracy of ChatGPT on MMLU is 85.23% when $S_{c}=4/4$, but it dwindles to 40.93% at $S_{c}=3/4$ and further plummets to 7.59% at $S_{c}=2/4$. This observed trend, consistent across various settings, suggests that the consistency of predictions across re-ordered instances is indicative of the reliability of the predictions. Consequently, we posit the following hypothesis:
Proposition 3.1.
An LLM can be deemed proficient in answering a specific MCQA-format question only if it consistently predicts the same correct answer across all permutations of option orders.
Table 4: Consistency of LLMs between the MCQA-format and True/False-format questions. correct: True/False questions derived from the correct MCQA options. distracting: True/False questions derived from the distracting MCQA options. † denotes the experiments on the sub-datasets where the LLMs predict correctly with the original MCQA settings.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
True or False: ${}^{{\dagger}}$ 0-shot 70.31 70.49 88.20 60.98 87.62 76.58
5-shot 66.32 71.75 91.04 63.70 87.76 75.58
True or False: ${}^{{\dagger}}$ 0-shot 52.52 55.77 48.51 70.74 68.48 69.87
5-shot 55.43 55.56 51.75 79.39 68.88 71.62
Altered Option with “None of the Above”
Expanding on previous research conducted with private LLMs (Kadavath et al., 2022), we investigate the impact of introducing a “None of the above” option in place of the correct answer within the test datasets to assess the adaptability of LLaMA, Baichuan, and ChatGPT. As delineated in Table 3, this modification precipitates an extreme decline in performance for the LLMs, with a reduction in both accuracy (Acc) and accuracy-hard (Acc-H) across all models by as much as 70.9%, from 74.44% to a mere 3.54%. Notably, except in the 5-shot scenario for ChatGPT, the models invariably fail to select the “None of the above” option as a substitute for the correct answer, with Acc-H falling below 5% in all other scenarios, underscoring the deficiency in the LLM ability to truly discern the knowledge within the MCQA-format questions.
3.4 Transformation to True-or-False Format
Beyond the alterations in MCQA answer options, we transform the original MCQA-format tasks into a True-or-False format, as depicted in Figure 2. This transformation is exclusively applied to the instances where the LLMs have demonstrated consistent and correct performance (with an $S_{c}$ of $1$). For every MCQA instance with four answer choices, we generate one True-or-False-format question using the correct option and three additional questions using the distracting options, anticipating that the LLMs will respond accurately with “Yes” and “No” respectively. Table 4 elucidates the degree of consistency between the original MCQA-format questions and the adapted True-or-False-format ones. Notably, LLMs exhibit a higher propensity for affirming True-or-False questions that originate from previously selected correct options, evidenced by a consistency rate of up to 91.07%. Conversely, the models demonstrate a reluctance to negate True-or-False questions derived from distracting options. This is evidenced by the fact that up to 51.49% of all distracting options are predicted as correct in the True-or-False format, signifying a divergence in the model decision between MCQA and True-or-False-format questions and further implying the unreliability of MCQA benchmarks.
3.5 Summary
This section has presented a comprehensive experimental analysis, revealing the manifestation of REsponse VAriability Syndrome (REVAS) within Large Language Models. Our empirical findings cast a shadow on the reliability of MCQA as a measure of LLM capabilities, as evidenced by the inconsistent responses of LLMs to derivations of the same MCQA-format question. The identification of REVAS in LLMs is pivotal in recognizing the inherent limitations of current evaluation strategies. Afterward, we will delve into the causative factors for the observed behavior patterns.
4 The More Correct but not the Only Correct.
Figure 4: The propensity of LLMs to select the most correct choice from a set of options and elucidating experiments.
The preceding section has elucidated the enigmatic phenomenon of REVAS, characterized by the unreliable consistency in the responses of LLMs, notably perturbed by minor alterations in the answer options. This inconsistency becomes extreme with the incorporation of the “None of the above” option, often leading to poor performance in MCQA tasks. With these observations, we hypothesize the following:
Assumption 4.1.
As probabilistic generative constructs, LLMs are prone to opt for the most correct option in the MCQA-format task, as opposed to the solely correct option.
Substituting the correct option with “None of the above”, with the above assumption, steers the LLMs towards the selection of the second most correct option, as depicted in Figure 4. To substantiate our assumption, we elaborate experiments encompassing: (1) Option probability analysis; (2) MCQA incorporating multi-selections; and (3) MCQA with the substitution of misleading options, as illustrated in Figure 4. The latter duo of experiments scrutinizes instances where LLMs have erroneously predicted under the original MCQA scenarios.
Table 5: Confidence of specific options in MCQA tasks with LLaMA2-13B and Baichuan2-13B. $C_{c}$: mean confidence of the correct options. $C_{d}$: mean confidence of the most distracting options (with the highest confidence). $C_{d}/C_{c}$: relative confidence score of $C_{d}$ divided by $C_{c}$. † denotes the experiments on the sub-datasets where the LLMs predict correctly with the original MCQA settings.
MMLU MedMCQA
$\textbf{L}^{{\dagger}}$ $C_{c}$ 35.04 35.96
0-shot $C_{d}$ 32.48 30.15
$C_{d}/C_{c}$ 92.69% 83.84%
$C_{c}$ 34.53 35.60
5-shot $C_{d}$ 31.91 30.32
$C_{d}/C_{c}$ 92.41% 85.17%
$\textbf{B}^{{\dagger}}$ $C_{c}$ 65.90 66.32
0-shot $C_{d}$ 61.37 60.08
$C_{d}/C_{c}$ 93.13% 90.59%
$C_{c}$ 65.36 67.30
5-shot $C_{d}$ 61.13 62.12
$C_{d}/C_{c}$ 93.53% 92.30%
Option Probability Analysis
In instances where LLMs regard certain distracting options in the MCQA as “correct but not the most correct”, it is plausible for the LLMs to acknowledge these distracting options as correct in other contexts. The aforementioned experiments reveal that question format is a significant factor in model performance. Consequently, we solely consider instances manifesting absolute consistency in all derived instances from the original instance, imposing a more stringent constraint. We probe the confidence of LLMs, as denoted in Chen et al. 2023, by examining the distribution of token probabilities of answer option tokens within the vocabulary list, aiming to discern the existence of “correct” versus “more correct” phenomena. By reformulating MCQA questions into True-or-False questions with both correct and distracting answer options, as represented in Figure 4, we derive independent confidence scores for each option with the probability of token “yes” instead of that of symbol options. Table 5 demonstrates the mean confidence levels of LLMs on the correct ($C_{c}$) and the most distracting ($C_{d}$ the distractor with the highest confidence) options, along with the confidence scores $C_{d}/C_{c}$ to assess the confidence level of $C_{d}$ relatively. Experimental results show that although Baichuan exhibits higher confidence than LLaMA in questions with the correct option, both LLMs attribute a degree of correctness to the most distracting options, with confidence levels exceeding 90% of that associated with the correct options in the MMLU. Consequently, despite accurate predictions and consistency across re-ordered options, LLMs may still perceive certain distracting options as correct, though to a lesser extent compared to the ground-truth options.
Multi-selection in MCQA
Conventionally, LLMs in MCQA paradigms are mostly expected to identify a single correct option. If the most probable option of an LLM diverges from the correct option, erroneous predictions will ensue. Furthermore, as for the wrong cases, if LLMs are prompted to identify all correct options and the correct options are encapsulated within the responded selections, it substantiates the preference of LLMs for “more correct” options over exclusively correct ones. We prompt the LLMs for multiple potential answers, collating instances where the LLMs have made mistakes on the original MCQA tasks. Herein, we designate the distracting options, erroneously identified as correct by the LLMs, as misleading options. Table 6 showcases the recall rates of correct and misleading options in LLM responses under the multi-selection MCQA paradigm, alongside the proportion of instances where LLMs render multiple selections. Results manifest that, in spite of the variance in multi-selection MCQA performance across different LLMs, correct options are encapsulated within the selections in over 78% of instances (up to 92.5%), a metric that surpasses the recall rates for misleading options (only up to 70.31%) previously predicted by the LLMs. This intriguing inconsistency in the recall of correct and misleading options intimates that LLMs not only reclassify unchosen correct options as correct in the multi-selection paradigm but also reassess their stance on misleading options between the single-selection and multi-selection MCQA paradigms.
Table 6: Recall of the correct and misleading options in the responses of LLMs with the setting of multiple selections for MCQA under the few-shot setting. %multi: the proportion of instances where the LLMs generate multiple selections. $R_{correct}$: recall of the correct options in the multiple selections. $R_{misleading}$: recall of the misleading options in the multiple selections.
MMLU MedMCQA
LLaMA2-13B %multi 23.12% 24.38%
$R_{correct}$ 92.50 78.91
$R_{misleading}$ 67.50 70.31
Baichuan2-13B %multi 82.77% 77.68%
$R_{correct}$ 86.29 88.59
$R_{misleading}$ 52.28 48.22
ChatGPT 3.5 %multi 62.14% 68.81%
$R_{correct}$ 85.06 81.27
$R_{misleading}$ 66.67 69.66
MCQA with the Misleading Options Replacement
In line with our definition of misleading options mentioned above, we explore whether replacing these misleading options with arbitrary tokens directs LLMs toward the correct options. This also encompasses instances in which LLMs have misjudged under the original MCQA scenarios. Table 7 elucidates that the LLMs rectify their initial misjudgments in from 39.23% to 57.99% of all instances, evidencing the influence of misleading options on their predictions. These findings also insinuate that certain instances may be intrinsically challenging for LLMs, and substituting a single misleading option may only marginally enhance model performance. The fundamental lack of pertinent knowledge may result in the erroneous predictions of the LLMs.
Table 7: Ratio of the instances where the LLMs turn to predict correctly with the replacement of misleading options under the few-shot setting.
MMLU MedMCQA
LLaMA2-13B 57.99 40.96
Baichuan2-13B 48.55 39.23
ChatGPT 3.5 53.57 44.47
5 LLMs have mostly Learned from the Right
During the stages of pre-training and supervised fine-tuning (SFT), GPT-like LLMs undergo training to predict subsequent text tokens by optimizing the cross-entropy loss function with massive texts. This paradigm predominantly concentrates on the “correct” sequences within textual contexts, given that the bulk of knowledge is mostly encapsulated in assertive statements. This approach, however, may inadvertently neglect a comprehensive understanding of absolute “correctness”, that alternatives to the “correct” answer are inherently incorrect. In light of this observation, we propose the ensuing assumption:
Assumption 5.1.
LLMs predominantly derive knowledge from “correct” contexts encountered during pre-training and supervised fine-tuning, potentially neglecting the learning from “incorrect” contexts.
To empirically evaluate this hypothesis, we employ the Baichuan2-7B-Base model, which is solely pre-trained, and utilize the MedMCQA dataset to curate distinct training and testing sets. Specifically, we extract 10k and 2k non-overlapping instances from MedMCQA, designated as $M_{10k}$ and $M_{2k}$ respectively. The training set comprises: (1) 10k “positive” instances $T_{Pos}$, formulated from questions paired with the correct answers, and (2) 10k “negative” instances $T_{Neg}$, assembled from questions coupled with distracting options. Both subsets, $T_{Pos}$ and $T_{Neg}$, are derived from $M_{10k}$ for the purpose of pre-training. Subsequently, we assemble 2k out-of-distribution instances $T_{SFT}$, extracted from $M_{2k}$, to facilitate the LLMs to learn across diverse question formats during supervised fine-tuning. For the testing phase, we employ an integrated task comprising a balanced amalgamation of original MCQA-format questions, modified MCQA-format questions incorporating the option “None of the above”, and True-or-False-format questions, all sourced from the $M_{10k}$ instances. These testing instances necessitate the same foundational knowledge as those within the $T_{Pos}$ and $T_{Neg}$ sets, albeit presented in varied question formats.
Employing the base model, we execute three distinct training strategies: (1) Only SFT on $T_{SFT}$ to establish a baseline performance. (2) Continuation of pre-training with positive instances $T_{Pos}$, followed by SFT on $T_{SFT}$, to analyse the influence of positive instances. (3) A sequential training involving positive, negative, and SFT instances to ascertain the contribution of negative instances in enhancing model performance. Consistent training steps are maintained across the latter two strategies. The results, as tabulated in Table 8, display that while pre-training only with “positive” instances diminishes performance in both MCQA and True-or-False-format tests, the incorporation of “negative” instances induces a substantial enhancement, particularly in the True-or-False-format test, marked by a 30.53% improvement. This outcome underscores the pivotal role of “negative” instances in fostering a more comprehensive and nuanced learning process within LLMs.
Table 8: Schematic experiments of training the LLMs with positive, negative, and SFT instances on Baichuan2-7B-Base.
Training Strategy MCQA Test True/False Test
SFT 59.81 49.84
Pos + SFT 44.22 (-15.59) 33.01 (-16.83)
Pos + Neg + SFT 68.19 (+8.38) 80.37 (+30.53)
Impact Statements
The core objective of this research is to critically analyze and challenge the rationality of multiple-choice question-answering (MCQA) as a trustworthy evaluation task for Large Language Models (LLMs). This paper posits that the LLMs in the current era fail to maintain consistency in their responses across various derivations of the same question, thereby casting doubt on the authenticity of MCQA as a robust evaluation strategy. The implications of this study extend beyond academic discourse, potentially eliciting a paradigm shift in the automatic evaluation of LLMs.
Appendix A Details of the Performance of the LLMs in Figure 3.
Table 9: Performance of LLMs on the original and modified MCQA-format questions. Acc: accuracy of LLMs with the original order of answer options. $\bar{S}_{c}$: mean value of $S_{c}$. Acc-H: accuracy-hard, that the prediction of an instance is correct only when LLMs predict consistently across all re-ordered options of the instance.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
Original 0-shot Acc 72.73 43.23 70.21 45.81 65.00 56.88
$\bar{S}_{c}$ 75.06 68.12 83.13 75.75 84.79 79.71
Acc-H 26.99 12.82 39.38 21.20 44.69 34.44
5-shot Acc 74.44 43.33 71.26 43.95 89.80 59.77
$\bar{S}_{c}$ 79.70 67.55 70.08 62.05 83.68 78.92
Acc-H 30.97 12.93 29.64 13.96 43.36 31.97
Increase 6 0-shot Acc 43.30 35.88 74.71 44.47 74.03 56.46
$\bar{S}_{c}$ 25.66 25.90 74.00 65.33 78.98 72.16
Acc-H 0.00 0.10 29.42 14.37 38.72 27.51
5-shot Acc 50.57 43.43 41.03 37.23 57.79 57.00
$\bar{S}_{c}$ 74.50 63.11 62.28 55.09 71.12 72.90
Acc-H 22.35 8.79 19.91 9.62 30.71 34.51
Increase 8 0-shot Acc 28.76 33.71 44.44 44.26 83.33 58.01
$\bar{S}_{c}$ 25.06 25.16 72.07 61.40 76.49 69.75
Acc-H 0.00 0.00 26.33 12.31 36.06 24.82
5-shot Acc 77.22 43.23 62.22 41.88 75.79 57.08
$\bar{S}_{c}$ 74.17 63.57 67.64 56.39 72.90 67.45
Acc-H 28.32 10.55 22.79 10.86 34.51 25.65
Increase 10 0-shot Acc 49.06 33.30 75.45 44.57 70.00 56.77
$\bar{S}_{c}$ 25.00 25.05 68.36 56.15 74.06 66.93
Acc-H 0.00 0.00 21.90 11.07 29.20 22.65
5-shot Acc 75.29 42.81 76.09 41.78 84.44 60.19
$\bar{S}_{c}$ 74.17 58.07 64.99 54.27 69.91 67.79
Acc-H 28.98 9.41 20.13 10.86 30.97 26.89
Reduction 3 0-shot Acc 43.79 46.74 47.06 49.33 66.99 63.39
$\bar{S}_{c}$ 75.33 77.59 86.55 81.41 90.38 87.38
Acc-H 20.13 18.82 37.17 24.40 48.23 37.54
5-shot Acc 60.78 48.50 54.08 52.22 64.44 64.74
$\bar{S}_{c}$ 88.88 78.96 87.17 80.95 90.15 86.76
Acc-H 41.81 22.75 38.94 22.96 46.24 38.57
Reduction 2 0-shot Acc 48.72 50.88 60.26 59.88 87.95 68.98
$\bar{S}_{c}$ 84.79 84.77 93.36 90.18 95.58 93.51
Acc-H 32.52 29.78 59.51 44.26 43.58 37.13
5-shot Acc 62.75 51.60 60.13 61.53 85.71 74.46
$\bar{S}_{c}$ 93.58 84.64 93.25 90.07 95.13 93.15
Acc-H 62.39 29.27 59.96 43.64 41.59 35.88