### v2

# Beyond the Answers: Reviewing the Rationality of Multiple Choice Question Answering for the Evaluation of Large Language Models

Abstract
In the field of natural language processing (NLP), Large Language Models (LLMs) have precipitated a paradigm shift, markedly enhancing performance in natural language generation tasks. Despite these advancements, the comprehensive evaluation of LLMs remains an inevitable challenge for the community. Recently, the utilization of Multiple Choice Question Answering (MCQA) as a benchmark for LLMs has gained considerable traction. This study first investigates the limitations of MCQA as an evaluation method for LLMs and then analyzes the fundamental reason for the limitations of MCQA, that while LLMs may select the correct answers, it is possible that they also recognize other wrong options as correct. Finally, we propose a dataset augmenting method for Multiple-Choice Questions (MCQs), MCQA+, that can more accurately reflect the performance of the model, which underscores the need for more robust evaluation mechanisms in assessing the performance of LLMs.
1 Introduction
The emergence of Large Language Models (LLMs), such as GPT-3 [gpt3], LLaMA [touvron2023llama], and ChatGPT [chatgpt], represents a paradigm shift in the field of Natural Language Processing (NLP). These models have exhibited exceptional proficiency in mimicking human-like textual outputs, establishing their significance across a diverse array of applications. However, the challenge of effectively evaluating LLMs persists [chang2023survey]. This difficulty arises from the intricate nature of natural language. Conventional evaluation metrics for generative tasks often fall short in accurately assessing the performance of LLMs, since most LLMs can generate text contextually rich and coherent [thoppilan2022lamda], complicating the assessment of the outputs through merely quantitative evaluation based on text matching such as BLEU [papineni-etal-2002-bleu] and ROUGE [lin-2004-rouge].
Multiple-Choice Question Answering (MCQA) is a fundamental format for various tasks in NLP, such as common sense reasoning [csqa, siqa, swag], reading comprehension [race, cqa] and cloze-style tasks [hs, sc]. Each MCQA instance comprises a question paired with several answer options, requiring models to identify the correct response as depicted in Figure 1. As a non-subjective metric, MCQA serves as a prominent automatic evaluation method with accuracy as an evaluation metric for numerous LLMs to test for the commonsense knowledge or knowledge for specific domain [leo_gao_2021_5371629, touvron2023llama, openai2023gpt4].
Figure 1: An example of MCQA and two ranking strategies.
Despite the advanced performance of LLMs on the accuracy of MCQA-format benchmarks like MMLU [mmlu], previous studies have discussed a key challenge that persists in evaluating LLMs is maintaining invariability in responses when confronted with different orders of answer choices for a same question [robinson2022leveraging, wang2023large, zheng2023large], which underscores an issue that the accuracy of MCQA-format tasks may not reflect the authentic capability of LLMs. However, the above phenomenon may not be the only issue in the evaluation of LLMs with MCQA-format questions.
In this study, we conduct a detailed exploration through a series of experiments utilizing various configurations derived from original MCQA-format datasets. We aim to deepen the understanding of the inherent limitations of MCQA tasks as tools for evaluating LLMs with a focus on response invariability. Afterward, we explore the underlying reasons for the observed behavioral patterns in LLMs through substantial experiments, that LLMs probably choose the most correct from the answer options and meanwhile regard some other options as correct as well. Thus, evaluating LLM performance based on MCQA can yield ambiguous results. To address these challenges, we introduce an augmentation of the MCQA dataset, termed MCQA+, which includes variations of the original MCQs and is designed to more accurately reflect the true capabilities of LLMs. Empirical results demonstrate that the performance of LLMs on the MCQA+ dataset is significantly inferior to that on the original MCQA dataset, suggesting that our MCQA+ can facilitate the development of more robust and adaptable NLP models. This, in turn, may help narrow the gap between machine and human-like understanding and reasoning in NLP tasks.
In summary, our contributions are as follows:
  • We challenge the rationality of using MCQA benchmarks for evaluating LLMs by substantial experiments.
  • We elucidate that within the MCQA context, LLMs may recognize multiple options as correct, but they ultimately select the option they assess as the most accurate.
  • We propose a dataset augmentation method that expands the original MCQA into MCQA+, which better reveals the models’ capacities.
Table 1: Common LLMs and corresponding evaluation tasks involving MCQA.
Chinchilla Falcon FLAN Galactica GPT4 LLaMA2 Mistral PaLM2
MMLU
Big Bench
CommonsenseQA
AGI Eval
2 Limitations of MCQA Task as Benchmark for LLMs
2.1 MCQA Tasks as Benchmarks for LLMs
The assessment of LLMs has increasingly employed MCQA as a standardized metric. This method, featuring a single question accompanied by multiple answer choices, serves to gauge the performance of an LLM performance in specific tasks or domains. Notable MCQA datasets utilized in evaluating LLMs encompass a diverse range including multitask language comprehension (e.g., MMLU [mmlu]), benchmark assessments for LLMs (e.g., Big Bench [srivastava2023beyond]), commonsense reasoning (e.g., HellaSwag [hs], ARC [arc], PIQA [piqa], WinoGrande [w]), and human-centric evaluations (e.g., AGI-eval [zhong2023agieval]). Table 1 lists prominent LLMs such as Chinchilla [chinchilla], Falcon [almazrouei2023falcon], FLAN [chung2022scaling], Galactica [taylor2022galactica], GPT4 [openai2023gpt4], LLaMA2 [touvron2023llama2], Mistral [jiang2023mistral], and PaLM2 [anil2023palm2], along with their respective evaluations with MCQA-format tasks. Additionally, MCQA tasks extend to tests like AP exams for GPT4 and medical-specific MultiMedQA for Med-PaLM [singhal2023large]. The adoption of MCQA tasks as an evaluation of LLMs is based on the hypothesis that MCQA can effectively and accurately reveal the true capabilities of the LLMs. If so, LLMs must demonstrate an understanding of the contextual content and grasp the requisite knowledge for correctly answering questions, thereby exhibiting invariability across similar questions.
In this study, we select three widely recognized LLMs: LLaMA2 [touvron2023llama2], Baichuan2 [yang2023baichuan], and ChatGPT 3.5 [chatgpt], for evaluation on the MMLU [mmlu] benchmark and the MedMCQA [medmcqa] dataset. MMLU represents a prevalent benchmark in MCQA for LLMs as presented in Table 1, while MedMCQA demands extensive medical knowledge and is challenging for most LLMs. We adopt the greedy strategy during decoding to avoid the influence of token sampling. We conduct comprehensive experiments to demonstrate that the accuracy of LLMs on the MCQA tasks does not accurately reflect their capabilities concerning the specific knowledge required by the questions.
Figure 2: Experimental designs derived from original answer options in the format of MCQA.
2.2 Further Exploration with Re-ordered Answer Options
Prior research [robinson2022leveraging] has demonstrated that re-ordering answer options can lead to variability, where LLMs generate variant answers with the re-ordered answer options for the same question. In our study, we extend this exploration to include MMLU and MedMCQA datasets tested on LLaMA2, Baichuan2, and ChatGPT models. Our findings also corroborate the existence of such variability (details in Appendix A). Meanwhile, we prompt a deeper examination from two angles: (1) the impact of varying the number of answer options on performance variability; and (2) the correlation between this variability and the accuracy of model responses.
2.2.1 Variation in Answer Option Quantity
We explore the impact of varying the number of answer choices in MCQA tasks. Generally, MCQA-format tests feature four options per question [haladyna1993many]. We extend this framework by modifying the number of answer options to assess the adaptability of LLMs under these varied conditions.
When incrementing the number of answer options, we incorporate additional non-semantic tokens as distracting options to circumvent the unintentional introduction of plausible correct responses, thereby preserving the authenticity of the questions, as illustrated in Figure 2. This expands the number of answer options to 6, 8, and 10. In contrast, when reducing the number of options, we randomly eliminate distracting options, simplifying the questions by reducing the options to 3 or 2.
Figure 3: Performance of LLMs on MCQA tasks of MMLU and MedMCQA with variations in the number of options. “Accuracy” denotes the accuracy reported with the initial answer option order. “Acc-H” represents the proportion of datasets where LLMs answer correctly across all corresponding numbers of re-ordered instances of an original question. Marked values in the sub-figures have 4 original options under the 0-shot setting.
By combining the adaptions of re-ordering and quantity variation of options, Figure 3 demonstrates the performance of LLMs on the MMLU and MedMCQA tasks. The metrics “Accuracy” and “Acc-H” are the same defined as in the preceding section. Intuitively, with fewer answer options, LLMs are expected to exhibit improved accuracy (achieving better “Accuracy” and maintaining “Acc-H”) as we have reduced the number of distracting options. Conversely, with increased answer options, LLMs are anticipated to maintain their accuracy since only non-distracting tokens (non-semantic tokens) have been added.
However, taking LLaMA2-13B as an illustrative example, it exhibits an accuracy of 72.73% with the original set of four options on MMLU under the zero-shot setting. However, with 2 to 10 answer options (other than 4 options), a marked volatility in accuracy is noted, ranging from 28.76% to 49.06%. Meanwhile, Acc-H exhibits a general declining trend as the number of options increases. This pattern of fluctuating accuracy and diminishing Acc-H values is consistent across all three LLMs on the MMLU. In experiments involving MedMCQA, a similar downward trend is observed for Accuracy, and Acc-H with the increment in answer options. Remarkably, for LLaMA2-13B, Acc-H plummets to 0.00% with six or more answer options, casting doubts on the reliability of the LLMs with a larger array of answer options. These phenomena underscore the unreliability of the MCQA task when the number of answer options is altered, as the following observation:
Observation 1: LLMs exhibit “overfitting” to MCQs with exactly four options. Consequently, the performance on MCQs with fewer or more than four options is not readily extrapolatable.
2.2.2 Correlation between the Variability and Accuracy
First, we conduct experiments with MCQA in its original answer option, recording baseline accuracy as “Acc”. Second, for simplicity, we expand the answer options in each MCQA-format question into four random distinct orders to assess the variability of LLMs to option sequencing as an example shown in Figure 2. We propose a novel evaluation metric, termed accuracy-hard (“Acc-H”), which quantifies the percentage of instances where LLMs invariantly and correctly predict across all four re-ordered orders. For example, Table 2 demonstrates that although LLMs like LLaMA2 can achieve an accuracy of 72.73% on the original MMLU, the Acc-H dropping to 26.99% indicates that LLaMA generates many different results if provided with answer options in different answer option orders. Afterward, Table 2 further analyzes the invariability (consistency) of LLMs when faced with four re-ordered answer options under the zero-shot setting. “$c=3/4$” indicates the subset of instances where three out of four predictions generated by LLMs are identical. It is noticeable that most of the LLMs can show invariability of no less than $3/4$. Notably, a higher $c$ value correlates with higher accuracy in the LLM responses. For instance, the accuracy of ChatGPT on MMLU is 85.23% when $c=4/4$, but it dwindles to 40.93% at $c=3/4$ and further plummets to 7.59% at $c=2/4$. This observed trend, consistent across various settings, suggests that the invariability of model responses across re-ordered instances is indicative of the model accuracy, which can be formalized as follows:
Observation 2: When an LLM consistently selects the same response to an MCQ, irrespective of the order in the sequence of the answer options, there is a high probability that the selected answer is correct.
Furthermore, the demonstration of invariability by LLMs in specific MCQs prompts us to ask the following question:
Question 1: If an LLM exhibits invariability in certain MCQs, can it be inferred to have genuinely mastered the knowledge involved in the questions?
Table 2: Model variability on MMLU and MedMCQA datasets. “Acc” denotes the accuracy on the original dataset. “Acc-H” represents the proportion of datasets where LLMs answer correctly across all four re-ordered instances of an original question. Distribution of consistency (and corresponding accuracy) with the original answer options. “%$c=n/4$” indicates the percentage of instances where LLMs consistently predicted $n$ out of 4 questions with re-ordered options. N/A: not applicable. All experiments are under the 0-shot setting.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
Acc 72.73 43.23 70.21 45.81 65.00 56.88
Acc-H 26.99 12.82 39.38 21.20 44.69 34.44
$c=4/4$ 43.02 (62.74) 19.03 (67.39) 48.23 (81.65) 34.54 (61.38) 52.43 (85.23) 40.43 (85.17)
$c=3/4$ 32.69 (73.21) 37.83 (64.33) 38.27 (44.51) 40.74 (42.13) 35.17 (40.93) 39.81 (43.48)
$c=2/4$ 19.52 (23.17) 29.37 (56.52) 11.28 (35.29) 17.89 (31.79) 11.50 (7.59) 17.89 (14.07)
$c=1/4$ 4.77 (N/A) 13.77 (N/A) 2.22 (N/A) 6.83 (N/A) 0.90 (N/A) 1.87 (N/A)
3 Does Model Invariability Indicate True Knowledge Mastery?
Question 1 offers a potential perspective for evaluating LLMs using MCQA, suggesting that the results on data where the model exhibits invariability may, to some extent, be reliable. Nevertheless, subsequent experiments will challenge this question, and then raise a vital problem within the MCQA-based LLM evaluation through (1) substituting the correct answer options in the MCQs with “None of the above” options and (2) transforming the MCQs into True-or-False format.
3.1 Introducing a “None of the Above” Option
Building upon earlier studies with close-source LLMs [kadavath2022language], we explore the effect of substituting the correct answer option in the MCQs with a “None of the above” option, as illustrated in Figure 2. This modification is applied across the entire selected datasets of MMLU and MedMCQA, as well as on particular subsets (marked as $\blacklozenge$) where LLMs previously responded correctly to all re-ordered answer choices. If the answer to Question 1 is “yes”, in these subsets, LLMs are presumed to possess a genuine comprehension of the knowledge in the questions from the subsets.
As presented in Table 3, as for the experiments on the entire selected datasets, this substitution results in a dramatic decline in LLM performance, with a decrease in both accuracy (Acc) and accuracy-hard (Acc-H) metrics by as much as 70.9%, from 74.44% to a mere 3.54%. Notably, except in a 5-shot scenario involving ChatGPT, the models generally fail to select the “None of the above” option as a substitute instead of the correct answer, with Acc-H falling below 5% in all other scenarios. The performance degradation is even more pronounced within the subsets ($\blacklozenge$), where LLMs, excluding ChatGPT, exhibit even poorer results. Despite prior correct predictions in all re-ordered instances before introducing the “None of the above” option, the results indicate a “No” answer to Question 1, further emphasizing the inadequacies of LLMs in reliably discerning the knowledge encapsulated in MCQA-format questions.
Table 3: Performance deterioration in the LLMs post “None of the Above” substitution. “Acc” denotes the accuracy of LLMs with the initial order of the other options. “Acc-H” represents the proportion of datasets where LLMs answer correctly across all corresponding numbers of re-ordered instances of an original question. $\blacklozenge$ means the subsets of datasets where LLMs have answered correctly across all re-ordered answer options before the alteration of “None of the above” with a baseline accuracy of 100%. Red font delineates the performance decrement compared to that before substitution.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
0-shot Acc 16.59 (-56.14) 20.27 (-22.96) 11.73 (-59.48) 14.99 (-30.82) 15.93 (-49.07) 27.09 (-29.79)
Acc-H 0.22 (-26.77) 1.34 (-11.48) 3.32 (-26.06) 7.24 (-13.96) 4.65 (-40.04) 8.07 (-26.37)
5-shot Acc 3.54 (-70.90) 4.45 (-38.88) 11.73 (-59.53) 14.99 (-28.96) 34.73 (-55.07) 40.85 (-18.92)
Acc-H 0.44 (-30.53) 0.31 (-12.62) 1.11 (-28.53) 1.76 (-12.20) 9.29 (-34.07) 15.51 (-16.46)
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
0-shot Acc 17.17 (-82.83) 12.28 (-87.72) 14.40 (-85.60) 10.32 (-89.68) 23.72 (-76.28) 36.04 (-63.96)
Acc-H 1.01 (-99.99) 0.00 (-100.00) 3.20 (-96.80) 5.56 (-94.44) 8.91 (-91.09) 15.62 (-84.38)
5-shot Acc 6.72 (-93.28) 0.00 (-100.00) 11.72 (-88.28) 4.55 (-95.45) 48.47 (-51.53) 41.67 (-58.33)
Acc-H 0.84 (-99.16) 0.00 (-100.00) 2.34 (-97.66) 0.00 (-100.00) 15.82 (-84.18) 16.67 (-83.33)
3.2 Transforming to True-or-False Format
Next, we transform the original MCQA-format tasks into a True-or-False format. For every MCQA instance with four answer choices, we generate four True-or-False-format questions, including one with the right option (T/F: right) and three questions with the wrong options (T/F: wrong) as depicted in Figure 4, anticipating that the LLMs will respond accurately with “Yes” and “No” respectively. This transformation is exclusively applied to the subsets where the LLMs have demonstrated invariant and correct performance (denoted as $\blacklozenge$).
Figure 4: True-or-false questions derived from the MCQs.
Table 4 provides an analysis of LLM performance on the True-or-False-format questions. Should Question 1 be true, LLMs are expected to exhibit near-perfect accuracy in this context. In practice, while LLMs can achieve an accuracy of up to 91.07% on the T/F: right datasets, the models demonstrate a reluctance to negate statements in the True-or-False questions derived from wrong options. This is evidenced by the fact that up to 51.49% of T/F: wrong datasets are predicted as right in the True-or-False format, signifying a divergence in the model decision between MCQA and True-or-False-format questions and implying the unreliability of MCQA benchmarks.
Table 4: Performance of LLMs on the True-or-False questions. $\blacklozenge$ means the subsets of datasets where LLMs have answered correctly across all re-ordered answer options of the MCQs.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
T/F: right 0-shot 70.31 70.49 88.20 60.98 87.62 76.58
5-shot 66.32 71.75 91.04 63.70 87.76 75.58
T/F: wrong 0-shot 52.52 55.77 48.51 70.74 68.48 69.87
5-shot 55.43 55.56 51.75 79.39 68.88 71.62
The experiments conducted have yielded the following insights:
Observation 3: While LLMs can display invariability on specific MCQs, their performance noticeably declines when confronted with modified MCQs under the following scenarios: (1) the correct answer options are replaced by “None of the above” and (2) dealing with True-or-False questions formulated from incorrect answer options.
3.3 The Most Correct One but not the Only Correct One.
Observation 3 suggests that LLMs may generate unexpected predictions when presented with modified questions lacking the original right answer option from the corresponding MCQs. Consequently, we can formulate the following intuitive Assumption 1, as illustrated in Figure 5. To validate Assumption 1, we elaborate a series of experiments.
Assumption 1: While LLMs display invariability on specific MCQs with a consistent answer option, they may regard this option as the most accurate, but not exclusively so. Specifically, LLMs might acknowledge the potential correctness of other wrong options, though not more correct than the chosen options.
Figure 5: Illustration of Assumption 1: LLMs may just select the most correct choice from an option set.
Figure 6: Experiments validating Assumption 1. “Option probability analysis” investigates the MCQs that LLMs consistently and accurately predict on all re-ordered orders. “Multiple selections” and “Replacement of misleading options” focus on the MCQs that LLMs erroneously predict.
Option Probability Analysis
Utilizing MMLU and MedMCQA, we investigate the confidence of LLMs, as outlined in [chen-etal-2023-close], by analyzing the distribution of token probabilities for answer option tokens. By converting MCQs into True-or-False questions with both right and wrong answer options, as represented in Figure 6, we derive the confidence scores based on the token probability for each answer option. Table 5 demonstrates the mean confidence of LLMs for the right options ($\text{Conf}_{right}$) and the wrong options ($\text{Conf}_{wrong^{\ast}}$, the wrong option that scores with the highest confidence in all wrong options for each MCQ), along with the relative confidence scores. The experimental results show that while LLMs consistently consider the wrong options as less correct than the right options, demonstrated by all relative confidence being below 100%, the wrong options still achieve substantial confidence ranging from 83.84% to 92.69% of those for the right options. Consequently, despite accurate predictions and invariability across re-ordered options, LLMs may still perceive certain wrong options as correct, though to a lesser extent compared to the right ones.
Table 5: Confidence of answer options in MCQA tasks with open-source LLMs. $\text{Conf}_{right}$ : mean confidence of the right options. $\text{Conf}_{wrong^{\ast}}$: mean confidence of the wrong options that score with the highest confidence. $\text{Conf}_{wrong^{\ast}}/\text{Conf}_{right}$: relative confidence score. $\blacklozenge$ denotes the experiments on the sub-datasets where the LLMs predict correctly with the original MCQA settings.
LLaMA2-13B Baichuan2-13B
MMLU MedMCQA MMLU MedMCQA
$\text{Conf}_{right}$ 35.04 35.96 65.90 66.32
0-shot $\text{Conf}_{wrong^{\ast}}$ 32.48 30.15 61.37 60.08
$\text{Conf}_{wrong^{\ast}}/\text{Conf}_{right}$ 92.69% 83.84% 93.13% 90.59%
$\text{Conf}_{right}$ 34.53 35.60 65.36 67.30
5-shot $\text{Conf}_{wrong^{\ast}}$ 31.91 30.32 61.13 62.12
$\text{Conf}_{wrong^{\ast}}/\text{Conf}_{right}$ 92.41% 85.17% 93.53% 92.30%
MCQA with Multiple Selections
For the MCQA datasets involved in this study, LLMs are tasked with identifying only one correct option per MCQ. We collect the instances where LLMs incorrectly predict the answers, denoted as MMLU and MedMCQA. In the above instances, the wrong options which LLMs have regarded as the right ones mistakenly are defined as misleading options. Then, LLMs are prompted to recognize all plausible correct options among all answer options. Table 6 showcases the recall for the right and misleading options, along with the proportion of instances where LLMs render multiple selections. The results reveal that the right options are included in the selections in over 78% of instances, reaching up to 92.5%. This indicates that the LLMs also recognize the right options as correct but less correct than the misleading ones.
[…]
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
Multi-selections %multi 23.12% 24.38% 82.77% 77.68% 62.14% 68.81%
Recallr 92.50 78.91 86.29 88.59 85.06 81.27
Recallm 67.50 70.31 52.28 48.22 66.67 69.66
Replacement Accuracy 57.99 40.96 48.55 39.23 53.57 44.47
MCQA with the Misleading Option Replacement
Apart from the multi-selection scenario, we explore the impact of replacing misleading options with arbitrary non-semantic tokens in MMLU and MedMCQA. Table 6 elucidates that the LLMs correctly identify the right options in 39.23% to 57.99% of instances, highlighting the influence of misleading options on their predictions. For cases where LLMs continue to make incorrect predictions, a fundamental deficit in relevant knowledge likely underpins the LLM incapability to generate the correct answers.
The analyses conducted across the three experimental scenarios support the likelihood of Assumption 1 being valid. This leads to an observation that underscores a fundamental limitation in using MCQA-based evaluations to accurately gauge the capabilities of LLMs:
Observation 4: In the context of MCQA, while LLMs may select the correct answer, there is a possibility that they also recognize other wrong options as correct.
4 MCQA+: Enhancing MCQA for Robust LLM Evaluation
Experimental analyses have revealed the limitations of the MCQA benchmark for evaluating LLMs. In response, we propose an augmenting approach based on the original MCQA dataset referring to empirical findings in the above observations, termed MCQA+. The MCQA+ dataset comprises (1) the original MCQs; (2) MCQs with re-ordered answer options; (3) MCQs featuring a different number of answer options; (4) MCQs where the correct options are replaced with “None of the above”; and (5) True-or-False questions derived from both right and wrong answer options. We convert each instance from the original MCQA format to the MCQA+ format and adopt the accuracy as the evaluation metric for LLMs (details provided in Appendix B).
Table 7 illustrates the comparative performance on the original MCQA dataset and the MCQA+ dataset. For LLaMA2 and Baichuan2, performance on the MMLU benchmark shows a substantial decline in both the 0-shot and 5-shot scenarios. For instance, accuracy for LLaMA2 plummeted from 72.73% to 45.21% in the 0-shot setting. This suggests that LLaMA2 and Baichuan2 might have been overfitted to the widely-used MMLU benchmark. For ChatGPT, a performance decrease is also evident under the 5-shot setting on MMLU. In contrast, performance on the MedMCQA dataset remains relatively stable across all three LLMs, possibly due to its specialized medical context. Notably, while LLaMA2 and Baichuan2 initially outperform ChatGPT on the original MMLU dataset, ChatGPT achieves markedly higher accuracy with MCQA+, suggesting its utility in revealing true model capabilities.
Although the MCQA+ evaluation provides a more reliable measure of LLM capabilities, it entails considerably higher computational costs compared to the original MCQA. To address this, we implement a sampling method for each MCQ instance, randomly selecting one from the expanded MCQA+ instances to form MCQA+ (×1)($\times$ 1)( × 1 ), which requires the same computational cost as the original MCQA. Table 7 confirms that even with this cost-efficient approach, MCQA+ (×1)($\times$ 1)( × 1 ) effectively exposes true capabilities.
Table 7: Model performance on the original MCQA dataset and MCQA+ with the accuracy metric.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
0-shot MCQA 72.73 43.23 70.21 45.81 65.00 56.88
MCQA+ 45.21 (-27.52) 41.20 (-2.02) 55.53 (-14.68) 47.52 (+1.71) 67.02 (+2.02) 56.50 (-0.38)
MCQA+ (×$\times 1$) 46.63 (-26.10) 40.07 (-3.16) 55.67 (-14.55) 44.71 (-1.10) 66.13 (+1.13) 56.00 (-0.88)
5-shot MCQA 74.44 43.33 71.26 43.95 89.80 59.77
MCQA+ 57.46 (-16.98) 42.37 (-0.96) 54.65 (-16.61) 45.13 (+1.18) 70.24 (-19.56) 61.04 (+1.27)
MCQA+ (×$\times 1$) 57.85 (-16.59) 40.88 (-2.45) 54.13 (-17.13) 45.89 (+1.94) 71.44 (-18.36) 61.03 (+1.26)
5 Related Work
Transformer-based language models [vaswani2017attention] have led the research of NLP into a new era. Prior work has elucidated the scaling law, positing that the efficacy of language models is typically proportional to both the model size and the magnitude of pre-training data [henighan2020scaling, kaplan2020scaling]. This principle has catalyzed a paradigm shift towards LLMs, with the number of parameters escalating from millions [bert, gpt2] to billions [radford2018improving, gpt3]. Recent advancements, including supervised fine-tuning and alignment with human values [instruct, chung2022scaling, bai2022training], have further augmented the capabilities of LLMs, enabling them to adhere more closely to human instructions and ethical considerations. Nonetheless, challenges persist since LLMs may show variability in the model responses, especially under the scenarios of MCQA. [robinson2022leveraging] termed the ability to associate the answer options and corresponding symbols as multiple choice symbol binding (MCSB) and proved that the MCSB ability varied significantly by models. Additionally, [wang2023large] revealed vulnerabilities in the ranking of candidate responses, which could be manipulated by altering the presentation order. [zheng2023large] investigated the token bias in LLMs, which could lead to selection biases during option prediction. Furthermore, [kadavath2022language] explored the reliability of the LLM performance and calibration, focusing exclusively on a set of private models under the MCQA settings. Their methodology involved modifying the options and transforming the original questions into binary-choice questions answering with options of “True” and “False”. Building on these foundational studies, in this study we conduct a comprehensive range of experiments applied to a diverse series of LLMs, including both open-source and commercial models. We aim to assess their performance invariability across various settings while using the same set of instances, thereby providing a broader understanding of their capabilities and limitations in different application scenarios.
6 Conclusion
In this study, we critically examined the efficacy of MCQA as a metric for evaluating LLMs, identifying key limitations of MCQA in capturing the true model capacities of LLMs. Our analysis revealed a tendency for overfitting in the context of four-option MCQs and highlighted potential behavioral logic issues that LLMs encounter with MCQs. The findings suggest that MCQA, in its conventional form, does not adequately reflect the full capacities of LLMs, as models may incorrectly recognize wrong options as correct. To overcome these deficiencies, we introduced MCQA+, an enhanced version of the original MCQA, designed to more accurately reveal the true capabilities of LLMs and improve the reliability of evaluation metrics.
[…]
Appendix A Variability of LLMs
Ideally, an intelligent LLM should display invariability to the re-ordered answer options for a specific MCQ, since the knowledge required by the question remains unchanged. We conduct experiments with MCQA in its original format on selected instances from MMLU and MedMCQA, recording baseline accuracy (“Acc”) just like the reported accuracy for evaluating LLMs. For a MCQ $q$, it includes four answer options $(opt_{1},opt_{2},opt_{3},opt_{4})$. We randomly permute the answer options into four orders, like:
$q_{1}=(q,A:opt_{1},B:opt_{2},C:opt_{3},D:opt_{4})$
$q_{2}=(q,A:opt_{2},B:opt_{4},C:opt_{1},D:opt_{3})$
$q_{3}=(q,A:opt_{2},B:opt_{1},C:opt_{3},D:opt_{4})$
$q_{4}=(q,A:opt_{4},B:opt_{1},C:opt_{3},D:opt_{2})$
LLMs are tested with the above questions with re-ordered answer options and generate answer options. Let $S=(opt_{i},opt_{j},\dots)$ be the set of answer options.
For $opt_{k}$ in $S$, $f(opt_{k})$ is the frequency of $opt_{k}$ in $S$. Subsequently, we introduce a consistency score “$S_{c}$” defined as follows:
$S_{c}=\frac{\max_{opt_{k}\in S}f(opt_{k})}{n}$ (1)
$\bar{S}_{c}=\frac{1}{N}\sum_{i=1}^{N}{S_{c}(i)}$ (2)
where $n$ represents the number of options in the MCQA task, and $N$ is the total number of task instances. For MCQA tasks with $n$ options, $S_{c}$ ranges from $1/n$ to $1.0$ describing the proportion of consistent answers with $n$ orders, where a score of 1.0 signifies consistent LLM performance across all re-ordered questions. Additionally, we propose a novel evaluation metric, termed accuracy-hard (“Acc-H”), which measures the proportion of instances where LLMs consistently and correctly predict across all re-ordered variations.
Apart from the analysis in Section 2.2.1, by combining the $S_{c}$, Figure 7 demonstrates a deeper analysis. The metric “Accuracy” is quantified based on questions with only a single option order for $n$ options. The $\bar{S}_{c}$ and Acc-H represent the mean values quantified across $n$ option orders for $n$ options respectively. Illustratively, taking Baichuan2-13B as an example, it exhibits an accuracy of 70.21% with the original set of four options on MMLU under the zero-shot setting, alongside a notable mean $S_{c}$ of 83.13. However, the Acc-H metric indicates a significant drop, with consistent and correct predictions in merely 39.38% of all instances when each instance is re-ordered into four distinct instances. Observing the performance of Baichuan2-13B with 2 to 10 answer options, a marked volatility in Accuracy is noted, ranging from 44.44% to 75.45%, with the peak accuracy surprisingly achieved with 10 options. Nonetheless, both $S_{c}$ and Acc-H exhibit a declining trend as the number of options increases. This pattern of fluctuating accuracy and diminishing $S_{c}$ and Acc-H values is consistent across all three LLMs on the MMLU. In experiments involving MedMCQA, a similar downward trend is observed for Accuracy, $S_{c}$, and Acc-H with the increment in answer options. Remarkably, for LLaMA2-13B, Acc-H plummets to 0.00% with six or more answer options, casting doubts on the reliability of the LLMs with a larger array of answer options. These observations underscore the questionable reliability of Accuracy metrics when based on a single order of answer options. Table 8 displays detailed results in Figure 7.
Figure 7: Performance of LLMs on MCQA tasks of MMLU and MedMCQA with a baseline of 4 (original) along with variations in the number of options. Accuracy is reported with the initial answer option order. Sc denotes the mean value of $S_{c}$. Acc-H: accuracy-hard, that the prediction of an instance is deemed correct only when LLMs predict consistently and correctly across all permutations of re-ordered options for a given instance. Marked values in the sub-figures are with 4 original options under the 0-shot setting.
Table 8: Performance of LLMs on the original and modified MCQA-format questions. Acc: accuracy of LLMs with the original order of answer options. $\bar{S}_{c}$: mean value of $S_{c}$. Acc-H: accuracy-hard, that the prediction of an instance is correct only when LLMs predict consistently across all re-ordered options of the instance.
LLaMA2-13B Baichuan2-13B ChatGPT 3.5
MMLU MedMCQA MMLU MedMCQA MMLU MedMCQA
Original 0-shot Acc 72.73 43.23 70.21 45.81 65.00 56.88
$\bar{S}_{c}$ 75.06 68.12 83.13 75.75 84.79 79.71
Acc-H 26.99 12.82 39.38 21.20 44.69 34.44
5-shot Acc 74.44 43.33 71.26 43.95 89.80 59.77
$\bar{S}_{c}$ 79.70 67.55 70.08 62.05 83.68 78.92
Acc-H 30.97 12.93 29.64 13.96 43.36 31.97
Increase 6 0-shot Acc 43.30 35.88 74.71 44.47 74.03 56.46
$\bar{S}_{c}$ 25.66 25.90 74.00 65.33 78.98 72.16
Acc-H 0.00 0.10 29.42 14.37 38.72 27.51
5-shot Acc 50.57 43.43 41.03 37.23 57.79 57.00
$\bar{S}_{c}$ 74.50 63.11 62.28 55.09 71.12 72.90
Acc-H 22.35 8.79 19.91 9.62 30.71 34.51
Increase 8 0-shot Acc 28.76 33.71 44.44 44.26 83.33 58.01
$\bar{S}_{c}$ 25.06 25.16 72.07 61.40 76.49 69.75
Acc-H 0.00 0.00 26.33 12.31 36.06 24.82
5-shot Acc 77.22 43.23 62.22 41.88 75.79 57.08
$\bar{S}_{c}$ 74.17 63.57 67.64 56.39 72.90 67.45
Acc-H 28.32 10.55 22.79 10.86 34.51 25.65
Increase 10 0-shot Acc 49.06 33.30 75.45 44.57 70.00 56.77
$\bar{S}_{c}$ 25.00 25.05 68.36 56.15 74.06 66.93
Acc-H 0.00 0.00 21.90 11.07 29.20 22.65
5-shot Acc 75.29 42.81 76.09 41.78 84.44 60.19
$\bar{S}_{c}$ 74.17 58.07 64.99 54.27 69.91 67.79
Acc-H 28.98 9.41 20.13 10.86 30.97 26.89
Reduction 3 0-shot Acc 43.79 46.74 47.06 49.33 66.99 63.39
$\bar{S}_{c}$ 75.33 77.59 86.55 81.41 90.38 87.38
Acc-H 20.13 18.82 37.17 24.40 48.23 37.54
5-shot Acc 60.78 48.50 54.08 52.22 64.44 64.74
$\bar{S}_{c}$ 88.88 78.96 87.17 80.95 90.15 86.76
Acc-H 41.81 22.75 38.94 22.96 46.24 38.57
Reduction 2 0-shot Acc 48.72 50.88 60.26 59.88 87.95 68.98
$\bar{S}_{c}$ 84.79 84.77 93.36 90.18 95.58 93.51
Acc-H 32.52 29.78 59.51 44.26 43.58 37.13
5-shot Acc 62.75 51.60 60.13 61.53 85.71 74.46
$\bar{S}_{c}$ 93.58 84.64 93.25 90.07 95.13 93.15
Acc-H 62.39 29.27 59.96 43.64 41.59 35.88
Appendix B Implementations
Construction of MCQA+
As described in Section 4, MCQA+ consists of 5 categories of instances derived from those in the original MCQA datasets. The number of instances in each category can be varied based on the specifics of the MCQA datasets. Our implementation of MCQA+ includes 14 times compared to the original MCQA dataset, encompassing the following variations:
  • One original MCQ;
  • Three MCQs with re-ordered answer options;
  • Five MCQs with varying numbers of answer options (2,3,6,8 and 10) as detailed in Section 2.2.1);
  • One MCQ where the correct answer options are replaced by “None of the above”;
  • Four True-or-False questions derived from the right and wrong answer options.
For the construction of MCQs in the MCQA+, a rule-based method is employed. And for the True-or-False questions, we utilize ChatGPT to convert the MCQs into True-or-False questions with prompts. Below is an example illustrating an original instance from MCQA and the augmented instances in MCQA+.
Figure 8: Example of instances in MCQA+ derived from the original MCQA.
Computational Resources
All experiments have been conducted on NVIDIA A800 PCIe.
Appendix C Limitations
Due to limited computational resources, this study was conducted using only the 13B parameter versions of LLaMA2 and Baichuan2, as well as ChatGPT3.5. The conclusions drawn are valid for these three models; however, it is possible that conclusions may differ from other models. From the data perspective, a subset of MMLU and MedMCQA was selected for experiments (complete selected data is provided in the supplemental materials). As this work focuses on analyzing the performance differences of LLMs across datasets derived from the same MCQA dataset, the test datasets, motivated by this focus, do not affect the conclusion that MCQA cannot precisely reflect the capabilities of LLMs. For the MCQA+, the computational cost of evaluation on the MCQA+ is multiple times more than that of the original MCQA dataset and we also provide a sampling approach considering the efficiency.
Appendix D Impact Statements
The core objective of this research is to critically analyze and challenge the rationality of multiple-choice question-answering (MCQA) as a trustworthy evaluation task for Large Language Models (LLMs). This paper posits that the LLMs in the current era fail to maintain consistency in their responses across various derivations of the same question, thereby casting doubt on the authenticity of MCQA as a robust evaluation strategy. The implications of this study extend beyond academic discourse, potentially eliciting a paradigm shift in the automatic evaluation of LLMs.